OpenStudy (anonymous):

Kendall scored 3 times as many points as Abdul scored in a basketball game. Otis scored 5 more points than Abdul scored. The three players scored 45 points all together. How many points did each player score?

OpenStudy (geneticrockhopper247):

If abdul scored 10 points, then kendall scored 30 points and otis scored 15. = 55 If abdul scored 7 points, then kendall scored 21 points and otis scored 12 = 40 If abdul scored 8 points, then kendall scored 24 points and otis scored 13 = 45 It's kinda a trial and error type thing. Start with an easy number like 10 and work your way up or down... Does that make sense?

OpenStudy (gorv):

let abdoul scored =x point

OpenStudy (gorv):

kendell=3*x

OpenStudy (gorv):

ottis=x+5

OpenStudy (gorv):

sum of their point = 45 x+3x+x+5=45

OpenStudy (gorv):

5x+5=45

OpenStudy (gorv):

subtract 5 from both side

OpenStudy (gorv):

5x=40

OpenStudy (gorv):

divide both side by 5

OpenStudy (gorv):

x=40/5=8

OpenStudy (gorv):

abdul=8

OpenStudy (gorv):

kendell=3*8

OpenStudy (gorv):

ottis=8+5